OpenStudy (anonymous):

The total number of students and adults at a concert was 460. The cost of tickets was $12 per student and $16 per adult. The total amount of income from tickets was $6520. How many students and how many adults were at the concert?

OpenStudy (anonymous):

Select the ordered pair from the choices below that is a solution to the following system of equations: 2y = 5x − 13 4x − 5y = 7 (-1/2,14/5) (3, 1) (-4, -1) (-3, -2)

OpenStudy (anonymous):

12x+16y=6520 x+y=460 where x=students y=adults

OpenStudy (anonymous):

You have a system of equations there are several ways to solve this. One way is through substitution. Do you know how to do that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

isolate one of the variables in equation 2 x+y=460 tell me what you get

OpenStudy (anonymous):

dude can you just solve it for me